OpenStudy (anonymous):

Someone please help!! Write the equation of a line ( y = mx + b ) with slope = -7/5 and with y-intercept = 6

OpenStudy (anonymous):

m is slope b is the y intercept

OpenStudy (anonymous):

let me show you the options

OpenStudy (anonymous):

y = (-7/5)x + 6 y = 6x - 7/5 y = (7/5)x - 6 None of the above

OpenStudy (anonymous):

( y = mx + b ) with slope = -7/5 and with y-intercept = 6 y=-7/5x+6

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

no problem;)